jhutter:

Katrina is building a deck. She cuts 3 lengths of wood from a 72-inch board. The first length is \(30\dfrac34\) inches. The second length is \(12\dfrac14\) inches. And the third length is \(16\dfrac14\) inches. How much of the original board remains after Katrina makes her cuts?

justus:

All you need to do is add them together then subtract 72 First step: Add the numbers 30 and 3/4+12 and 1/4+ 16 and 1/4=30+3/4+12+1/4+16+1/4 =30+12+16+3/4+1/4+1/4=58+1+1/4=59+1/4=59 and 1/4

justus:

Second step: Subtract from 72 72-59 and 1/4=72-(59+1/4)=72-59-1/4=71+1-59-1/4= 71-59+1-1/4= 71-59+4/4-1/4=12+3/4

justus:

The answer is 12 and 3/4 inches